OpenStudy (anonymous):

How do I write an expression that simplifies to 7, 13,or 17 by using the order of operations and at least one parentheses?

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

1(3 + 4) = 7 or 1(10 + 3) = 13 or 2 x 7 + (1 + 2) = 17

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

its ok....we are here to help...welcome to open study...have fun :)